OpenStudy (anonymous):

MEDAL AWARD. At most 500 ft of fencing is available for a steer pasture. There is only space for the pasture to be at most 100 ft wide. Write a system of two linear inequalities that describes this situation and write two possible solutions.

jimthompson5910 (jim_thompson5910):

Let's assume we have a rectangle Let x be the width and y be the length 

jimthompson5910 (jim_thompson5910):

It states that "At most 500 ft of fencing is available for a steer pasture" so the perimeter P is at most 500 ft, which means \[\Large P \le 500\]

jimthompson5910 (jim_thompson5910):

The perimeter of the rectangle above is P = 2(x+y), so... \[\Large P \le 500\] \[\Large 2(x+y) \le 500\] \[\Large x+y \le \frac{500}{2}\] \[\Large x+y \le 250\]

jimthompson5910 (jim_thompson5910):

"There is only space for the pasture to be at most 100 ft wide" means that \[\Large x \le 100\]

jimthompson5910 (jim_thompson5910):

And finally, the values of x and y cannot be negative (since they are the width and length) telling us that \[\Large x \ge 0\] \[\Large y \ge 0\]

jimthompson5910 (jim_thompson5910):

Put this all together and you get this system of inequalities \[\Large x+y \le 250 \\ \Large x \le 100 \\ \Large x \ge 0 \\ \Large y \ge 0\]

OpenStudy (anonymous):

I'm not following this.. There is only supposed to be 2 linear inequalities.. And when you solved p is less than or equal to 500, wouldn't you come out with the equation, y= -x+250?

jimthompson5910 (jim_thompson5910):

ok, so you can ignore the last two inequalities if you want and just focus on quadrant 1

jimthompson5910 (jim_thompson5910):

so you'd have the system \[\Large x+y \le 250 \\ \Large x \le 100\]

jimthompson5910 (jim_thompson5910):

And yes, you can solve for y to get \[\Large y \le -x+250\]

jimthompson5910 (jim_thompson5910):

the equation form of that is y = -x + 250 So you would graph the equation to form the boundary line